# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### 23 Dec 2023 v1

@@ 3.1. Depth-limited solving
Range $r_{i}\in\Delta\mathcal{S}_{i}(s_{0})$ for player $i$ in public state $s_{0}$ is a probability distribution over infosets $s_{i}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathcal{S}_{i}(s_{0})$ in public state $s_{0}$, which corresponds to player’s contribution to the probability of reaching this infoset if it followed the policy $\pi_{i}$. Furthermore range for public state $s_{0}$ is defined as $r=(r_{1},\dots,r_{N})\in\Delta\mathcal{S}_{1}(s_{0})\times\dots\times\Delta\mathcal{S}_{N}(s_{0})$. A public belief state is a public state with associated range $\beta=(s_{0},r)$. Optimal value function is a function $\boldsymbol{v}^{*}(\beta)\in\mathbb{R}^{|\mathcal{S}_{1}(s_{0})|\dots|\mathcal{S}_{N}(s_{0})|}$, that for given public belief state $\beta$ returns counterfactual values $v^{\pi}_{i}(s_{i})$ for each player $i$ and infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$ associated with public belief state $\beta$ if both players played optimally from that point onward based on associated ranges $r$. Value function $\boldsymbol{v}(\beta)$ is then defined similarly as optimal value function, but it’s output could be noisy. For any public belief state, we can construct subgame $\mathcal{G}(\beta)$, which is the same FOSG as $\mathcal{G}$, but it starts in $s_{0}$ with initial ranges $r$ instead of the initial state. Depth-limited subgame is defined as $\mathcal{G}^{\text{DL}}(\beta)$. However, after some depth limit, the value function $\boldsymbol{v}$ is called and returned counterfactual values are used as a terminal values.
@@
Resolving some subgame to improve blueprint policy $\pi$ naively would involve fixing policy $\pi$ for all players in previous parts of the game and using single $w^{A}$, where players act with corresponding ranges. However, this may result in a much more exploitable strategy Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017) than the original blueprint $\pi$. Multiple techniques were developed to deal with this issue in two-player zero-sum games Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017); Brown et al. (2020). In this work we will focus on the one proposed in Burch et al. (2014). When resolving from the perspective of player $i$, it constructs a slightly modified subgame called a gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$, where $v^{\pi}_{o}$ are counterfactual values for each infoset $s_{o}\in\mathcal{S}_{o}(s_{0})$ of opponent $o$ in a current public state $s_{0}$, when players play according to policy $\pi$. This subgame has the artificial state $w^{A}$ as an initial state, but instead of using full ranges $r$, it uses only ranges $r_{i}$ and $r_{c}$. Directly after this artificial state, $|\mathcal{H}(s_{0})|$ opponent’s decision nodes are separated into multiple infosets based on the information player has in a given history. The opponent $o$ has a choice to either terminate the game and receive the counterfactual value $v^{\pi}_{o}(s_{o})$ or continue the game as in the original subgame.
@@
Suppose that after the depth limit, the optimal value function is used and that the counterfactual values $v^{\pi^{{NE}}}_{o}(s_{o})$ used in the gadget game are computed for Nash equilibrium policy $\pi^{{NE}}$. Then the solution of gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$ is part of a Nash equilibrium in the entire game Burch et al. (2014); Kovařík et al. (2023).
@@
In this work, we focus on an alternative to the value function called multi-valued states Brown et al. (2018), which, after the depth limit, gives the opponent the choice to fix its strategy until the end of the game. Then, the game terminates with the reward corresponding to the reward if player $i$ follows some blueprint policy $\pi_{i}$, while the opponent follows the fixed strategy. Authors of Kovařík et al. (2023) pointed out that multi-valued states could be viewed as a form of a value function. We further discuss this in Section 4.2
3.2. Regularized Nash Dynamics
@@
Using usual reinforcement learning techniques to find an optimal policy with self-play does not guarantee convergence to Nash equilibrium and can be exploited Lanctot et al. (2017). However, regularizing the reward leads to establishing this guarantee Perolat et al. (2022); Sokota et al. (2023).
@@
$\mathcal{R}_{i}^{R}(h,a,\pi):=\mathcal{R}_{i}(h,a)-\eta\log(\frac{\pi(s_{i},a_{i})}{\pi^{R}(s_{i},a_{i})})+\eta\log(\frac{\pi(s_{-i},a_{-i})}{\pi^{R}(s_{-i},a_{-i})})$
@@
where $\pi^{R}$ is a regularization policy and $\eta$ is a hyperparameter for adjusting the regularization effect. Games with such a regularized reward have a single unique fixed point $\pi_{\text{fix}}$ for any regularization policy $\pi^{R}$ Perolat et al. (2021). Follow the Regularized Leader McMahan (2011) may be used to converge to $\pi_{\text{fix}}$. This $\pi_{\text{fix}}$ is not Nash equilibrium, but setting $\pi^{R}=\pi_{\text{fix}}$ will change the unique fixed point. If this is done repetitively, the fixed point will gradually converge to the Nash equilibrium. More regularizations were proposed in Sokota et al. (2023).
@@
In order to converge to a Nash equilibrium in large games with reward regularization, the authors in Perolat et al. (2022) propose using two-player V-trace with Neural Replicator Dynamics Hennes et al. (2020) as a loss function for policy head and usual L2 loss for value head. With this method, it is possible to achieve human-level performance in Stratego.
V-trace estimates infoset action values and infoset values with low bias and low variance for a policy $\pi$ by sampling full trajectories according to some sampling policy $\mu$. The estimate is done from the end of the trajectory to the start. Because of the importance sampling clipping, the estimated value is for a mix between $\pi$ and $\mu$, but the ratio may be tuned with a single hyperparameter Espeholt et al. (2018).
4. Multi-agent Search with Policy Transformation Value Network
@@
In this section, we present an algorithm Multi-agent Search with Policy Transformation Value Network (SePoT) that trains an additional critic alongside a policy network with some policy-gradient algorithm for imperfect information games, like RNaD Perolat et al. (2022). During gameplay, this critic is used to construct a depth-limited gadget game that is then solved with some search algorithm, in our case CFR+ Tammelin et al. (2015).
4.1. Network Training
@@
Performing a depth-limited search in imperfect information games requires a value function that, after the depth limit, assigns value to each infoset based on ranges Kovařík et al. (2023). In order to use a value function based on multi-valued states, it is necessary for each history $h$ to learn multiple values that correspond to the expected values if the opponent plays some fixed policies against the player’s blueprint policy. Training such value function presents two difficulties to overcome. The first is the selection of the opponent policies to be used as a basis to compute the expected values. In large games, it is not viable to simply use some predefined policy due to the sheer size that full policies have. The second is estimating the expected value of these policies.
@@
To find different opponent policies, we decided to use a similar method to the ”bias approach” in Brown et al. (2018). The main idea is to use just a single policy, which in our case is the policy from the policy network, and then applying some predefined mapping, that maps the network policy to a different policy. We call this mapping a policy transformation.
[…]
@@
Policy transformation $f^{T}:\Pi_{i}\times\mathcal{S}_{i}\rightarrow\Pi_{i}$ for player $i$ is a function that maps policy to a different one based on a current infoset.
@@
In poker, an example of a transformation is folding more frequently Bowling et al. (2009); Bard et al. (2014). Similarly, in Goofspiel, certain transformations may increase the probability of playing cards with lower value if the player is winning or decrease the probability of playing even-valued cards. These transformations do not have to be game-domain specific. However, the effectiveness of different transformations may vary across different games. These transformations aim to cover the strategy space as much as possible. Constant transformations that transform any policy into each pure strategy cover this space fully Brown et al. (2018); Kovařík et al. (2023). This work does not extensively delve into the specific methodology or criteria for selecting these transformations, and we leave it open for future work.
@@
$\begin{split}\mathcal{V}v_{i}(h)=&v_{i}(h)\\ +&\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}|h^{0}=h,\mu\right]\end{split}$ (1)
@@
{theorem}
(Inspired by Espeholt et al. (2018)): Let $\rho^{t}=\min{(\overline{\rho},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $c^{t}=\min{(\overline{c},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $\overline{\rho}\geq\overline{c}\geq 1$ and $\mu(s,a)>0\;\;\forall s,a$. Assuming there exists $\kappa\in(0,1]$, such that $\mathbb{E}_{\mu}\rho^{l}\Big{(}\prod_{{0\leq q<l}}c^{q}\Big{)}\geq\kappa$, where $l$ is a length of the trajectory. Then the V-trace operator $\mathcal{V}$ has a unique fixed point $v^{\pi_{\overline{\rho}}}$, which is the expected value of following policy
$\pi_{\overline{\rho}}(s_{i}(h^{t}),a_{i}^{t})=\frac{\min(\overline{\rho}\mu(s_{i}(h^{t}),a_{i}^{t}),\pi(s_{i}(h^{t}),a_{i}^{t}))}{\sum_{b\in\mathcal{A}_{i}(s_{i}(h^{t}))}\min(\overline{\rho}\mu(s_{i}(h^{t}),b),\pi(s_{i}(h^{t}),b))}$ (2)
@@
4.2. Constructing public belief state
@@
As discussed in the Section 2, naively employing search in the game could lead to highly exploitable strategies. Using resolving gadget Burch et al. (2014), which limits the opponent’s counterfactual value in root infosets, bounds the error after performing search in a subgame. Constructing a gadget game requires the opponent’s counterfactual values $v^{\pi}_{o}$ in each root infoset and public belief state $\beta=(s_{0},r)$. Most previous methods stored counterfactual values $v^{\pi}_{o}$, public state $s_{0}$ and ranges $r$ from previous searches Burch et al. (2014); Moravčík et al. (2017); Schmid et al. (2023). SePoT could leverage the same idea, but we propose a different approach, using the trained history network to compute counterfactual values, trained policy network to retrieve ranges for searching player $i$ and function that reconstructs the public state. This allows running the search from an arbitrary public state without the need to search in all parent public states.
@@
Safe search in imperfect information games has to be performed from all possible histories based on public information. This requires all histories $\mathcal{H}(s_{0})$ from current public state $s_{0}$ to construct the gadget game $\mathcal{G}^{G}_{i}$. Let us assume a function $f^{Rec}$ that generates all histories in the public state $s_{0}$ from the publicly available information to all players. While the details of reconstructing the public state $s_{0}$ are not explored in detail here, for games like Goofspiel or Battleships, the public state can potentially be reconstructed using Constraint Satisfaction Problem (CSP) Russell and Norvig (2009); Seitz et al. (2021), which we also used in our experiments.
@@
Gadget game $\mathcal{G}^{G}_{i}$ requires ranges for the player $i$ and the chance player. Chance player ranges $r_{c}$ can easily be stored during gameplay in the same way as older methods Moravčík et al. (2017); Schmid et al. (2023); Brown and Sandholm (2018). Suppose the player $i$ played only by using the policy from the neural network during current gameplay and decides to perform a search at some public state $s_{0}$. In that case, it has to recompute the ranges $r_{i}$ for each infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$. This requires backtracking from all of these infosets to all previous decision points and acquiring the policy from the policy network for each of them. Because we are in a perfect recall setting, each infoset cannot have more parenting infosets than the maximal trajectory length $l^{\text{max}}$ to this infoset from the initial state $w^{0}$, so this backtracking involves at most $|\mathcal{S}_{i}(s_{0})|l^{\text{max}}$ queries to the policy network.
@@
In order to compute the counterfactual values in root infosets for the opponent, let us assume that one of the transformations we used in training is identity transformation, meaning that the network output corresponds to the expected value if both players follow the policy from the network, which serves as an approximation of a Nash equilibrium. The opponent’s counterfactual values are the expected values multiplied by the reaches of both chance and resolving players. As discussed, these ranges are always available, so the counterfactual values can always be computed from the definition. Thanks to this, we have both counterfactual values $v^{\pi}_{o}$ and $\beta$ available at each point of the game, so we can construct a gadget game and perform a safe search.
4.3. Value Function

### 06 Feb 2024 v2

@@ 3.1 Depth-limited solving
Range $r_{i}\in\Delta\mathcal{S}_{i}(s_{0})$ for player $i$ in public state $s_{0}$ is a probability distribution over infosets $s_{i}~{}\in~{}\mathcal{S}_{i}(s_{0})$ in public state $s_{0}$, which corresponds to player’s contribution to the probability of reaching this infoset. Furthermore range for public state $s_{0}$ is defined as $r=(r_{1},\dots,r_{N})\in\Delta\mathcal{S}_{1}(s_{0})\times\dots\times\Delta\mathcal{S}_{N}(s_{0})$. A public belief state is a public state with associated range $\beta=(s_{0},r)$. Optimal value function is a function $\boldsymbol{v}^{*}(\beta)\in\mathbb{R}^{|\mathcal{S}_{1}(s_{0})|\dots|\mathcal{S}_{N}(s_{0})|}$, that for given public belief state $\beta$ returns counterfactual values $v^{\pi}_{i}(s_{i})$ for each player $i$ and infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$ associated with public belief state $\beta$ if both players played optimally from that point based on associated ranges $r$. Value function $\boldsymbol{v}(\beta)$ is then defined similarly, but it’s output could be noisy. For any public belief state, we can construct subgame $\mathcal{G}(\beta)$, which is the same FOSG as $\mathcal{G}$, but it starts in $s_{0}$ with initial ranges $r$ instead of the initial state. Depth-limited subgame is defined as $\mathcal{G}^{\text{DL}}(\beta)$. However, after some depth limit, the value function $\boldsymbol{v}$ is called and returned counterfactual values are used as a terminal values.
@@
Resolving some subgame to improve blueprint policy $\pi$ naively would involve fixing policy $\pi$ for all players in previous parts of the game and using single $w^{A}$, where players act with corresponding ranges. However, this may result in a much more exploitable strategy (Burch et al., 2014) than the original blueprint $\pi$. Multiple techniques were developed to deal with this issue in two-player zero-sum games (Burch et al., 2014; Moravcik et al., 2016; Brown and Sandholm, 2017; Brown et al., 2020). In this work we will focus on the one proposed in (Burch et al., 2014). When resolving from the perspective of player $i$, it constructs a slightly modified subgame called a gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$, where $v^{\pi}_{o}$ are counterfactual values for each infoset $s_{o}\in\mathcal{S}_{o}(s_{0})$ of opponent $o$ in a current public state $s_{0}$, when players play according to policy $\pi$. This subgame has the artificial state $w^{A}$ as an initial state, but instead of using full ranges $r$, it uses only ranges $r_{i}$ and $r_{c}$. Directly after this artificial state, $|\mathcal{H}(s_{0})|$ opponent’s decision nodes are separated into multiple infosets based on the information player has in a given history. The opponent $o$ has a choice to either terminate the game and receive the counterfactual value $v^{\pi}_{o}(s_{o})$ or continue the game as in the original subgame.
@@
Suppose that after the depth limit, the optimal value function is used and that the counterfactual values $v^{\pi^{{NE}}}_{o}(s_{o})$ used in the gadget game are computed for Nash equilibrium policy $\pi^{{NE}}$. Then the solution of gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$ is part of a Nash equilibrium in the entire game (Burch et al., 2014; Kovařík et al., 2023).
@@
In this work, we focus on an alternative to the value function called multi-valued states (Brown et al., 2018), which, after the depth limit, gives the opponent the choice to fix its strategy until the end of the game. Then, the game terminates with the reward corresponding to the reward if player $i$ follows some blueprint policy $\pi_{i}$, while the opponent follows the fixed strategy. Authors of (Kovařík et al., 2023) pointed out that multi-valued states could be viewed as a form of a value function. We further discuss this in Section 4.2
3.2 Regularized Nash Dynamics
@@
Using usual reinforcement learning techniques to find an optimal policy with self-play does not guarantee convergence to Nash equilibrium and can be exploited (Lanctot et al., 2017). However, regularizing the reward leads to establishing this guarantee (Perolat et al., 2022; Sokota et al., 2023).
@@
$\begin{split}&\mathcal{R}_{i}^{R}(h,a,\pi):=\\ &\mathcal{R}_{i}(h,a)-\eta\log(\frac{\pi(s_{i},a_{i})}{\pi^{R}(s_{i},a_{i})})+\eta\log(\frac{\pi(s_{-i},a_{-i})}{\pi^{R}(s_{-i},a_{-i})})\end{split}$ (1)
@@
where $\pi^{R}$ is a regularization policy and $\eta$ is a hyperparameter for adjusting the regularization effect. Games with such a regularized reward have a single unique fixed point $\pi_{\text{fix}}$ for any regularization policy $\pi^{R}$ (Perolat et al., 2021). Follow the Regularized Leader (McMahan, 2011) may be used to converge to $\pi_{\text{fix}}$. Setting $\pi^{R}=\pi_{\text{fix}}$ will change the unique fixed point. If this is done repetitively, the fixed point will gradually converge to the Nash equilibrium.
@@
In order to converge to a Nash equilibrium in large games with reward regularization, the authors in (Perolat et al., 2022) propose using two-player V-trace with Neural Replicator Dynamics (Hennes et al., 2020) as a loss function for policy head and usual L2 loss for value head. With this method, it is possible to achieve human-level performance in Stratego.
4 Multi-agent Search with Policy Transformations
@@
In this section, we present an algorithm Multi-agent Search with Policy Transformations (SePoT) that trains an additional critic alongside a policy network with some policy-gradient algorithm for imperfect information games, like RNaD (Perolat et al., 2022). During gameplay, this critic is used to construct a depth-limited gadget game that is then solved with some search algorithm, in our case CFR+ (Tammelin et al., 2015).
4.1 Network Training
@@
Performing a depth-limited search in imperfect information games requires a value function that, after the depth limit, assigns value to each infoset based on ranges (Kovařík et al., 2023). To use a value function based on multi-valued states, it is necessary for each history $h$ to learn multiple values that correspond to the expected values if the opponent plays some fixed policies against the player’s blueprint policy. Training such value function presents two difficulties to overcome. The first is the selection of the opponent policies to be used as a basis to compute the expected values. In large games, simply using some predefined policy is not viable due to the sheer size of full policies. The second is estimating the expected value of these policies.
@@
To find different opponent policies, we decided to use a similar method to the ”bias approach” in (Brown et al., 2018). The main idea is to use a single policy, which in our case is the policy from the policy network, and then apply some mapping that maps the network policy to a different policy. We call this mapping a policy transformation.
Definition 1.
@@
Policy transformation $f^{T}:\Pi_{i}\rightarrow\Pi_{i}$ for player $i$ is a function that maps policy to a different one.
@@
In poker, an example of a transformation is folding more frequently (Bowling et al., 2009; Bard et al., 2014). Similarly, in Goofspiel, certain transformations increase the probability of playing cards with lower value if the player is winning or decrease the probability of playing even-valued cards. These transformations do not have to be game-domain specific. However, the effectiveness of different transformations varies across different games. These transformations aim to cover the strategy space as much as possible. Constant transformations that transform any policy into each pure strategy cover this space fully (Brown et al., 2018).
In our experiments, we use transformations that move the actor policy $\pi$ in some direction with a step size $k$, where $k\in\mathbb{R}$ is a constant. These transformations are encoded within a neural network, which undergoes concurrent training with the actor network. At each training step, for each trajectory, we compute the direction in which the actor policy for that trajectory has shifted. Subsequently, we update only the transformation associated with the minimum Euclidean distance across the trajectory. Since each update should improve the actor policy against the current opponent’s policy, this direction should aim into the important regions of strategy space.
@@
$\begin{split}&\mathcal{V}v_{i}(h)=v_{i}(h)+\\ &\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}|h^{0}=h,\mu\right]\end{split}$ (2)
@@
Theorem 1.
(Inspired by (Espeholt et al., 2018)): Let $\rho^{t}=\min{(\overline{\rho},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $c^{t}=\min{(\overline{c},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $\overline{\rho}\geq\overline{c}\geq 1$ and $\mu(s,a)>0\;\;\forall s,a$. Let us assume there exists $\kappa\in(0,1]$, such that $\mathbb{E}_{\mu}\rho^{l}\Big{(}\prod_{{0\leq q<l}}c^{q}\Big{)}\geq\kappa$, where $l$ is a length of the trajectory. Then the V-trace operator $\mathcal{V}$ has a unique fixed point $v^{\pi_{\overline{\rho}}}$, which is the expected value of following policy
$\begin{split}\pi_{\overline{\rho}}(&s_{i}(h^{t}),a_{i}^{t})=\\ &\frac{\min(\overline{\rho}\mu(s_{i}(h^{t}),a_{i}^{t}),\pi(s_{i}(h^{t}),a_{i}^{t}))}{\sum_{b\in\mathcal{A}_{i}(s_{i}(h^{t}))}\min(\overline{\rho}\mu(s_{i}(h^{t}),b),\pi(s_{i}(h^{t}),b))}\end{split}$ (3)
@@
4.2 Constructing public belief state
@@
As discussed in the Section 2, naively employing search in the game could lead to highly exploitable strategies. Using resolving gadget (Burch et al., 2014), which limits the opponent’s counterfactual value in root infosets, bounds the error after performing search in a subgame. Constructing a gadget game requires the opponent’s counterfactual values $v^{\pi}_{o}$ in each root infoset and public belief state $\beta=(s_{0},r)$. Most previous methods stored counterfactual values $v^{\pi}_{o}$, public state $s_{0}$ and ranges $r$ from previous searches (Moravčík et al., 2017; Schmid et al., 2023). SePoT could leverage the same idea. Still, we propose a novel approach of using the trained history network to compute counterfactual values, the trained policy network to retrieve ranges for searching player $i$, and a function that reconstructs the public state. This allows running the search from an arbitrary public state without the need to search in all parent public states.
@@
Safe search in imperfect information games has to be performed from all possible histories based on public information. This requires all histories $\mathcal{H}(s_{0})$ from current public state $s_{0}$ to construct the gadget game $\mathcal{G}^{G}_{i}$. Let us assume a function that generates all histories in the public state $s_{0}$ from the publicly available information to all players. While the details of reconstructing the public state $s_{0}$ are not explored in detail here, for games like Goofspiel or Battleships, the public state can potentially be reconstructed using Constraint Satisfaction Problem (CSP) (Russell and Norvig, 2009; Seitz et al., 2021).
@@
Gadget game $\mathcal{G}^{G}_{i}$ requires ranges for the player $i$ and the chance player. Chance player ranges $r_{c}$ can easily be stored during gameplay like older methods (Moravčík et al., 2017; Schmid et al., 2023; Brown and Sandholm, 2018). Suppose the player $i$ played only by using the policy from the neural network during current gameplay and decides to perform a search at some public state $s_{0}$. In that case, it has to recompute the ranges $r_{i}$ for each infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$. This requires backtracking from all of these infosets to all previous decision points and acquiring the policy from the policy network for each of them. Because we are in a perfect recall setting, each infoset cannot have more parenting infosets than the maximal trajectory length $l^{\text{max}}$ to this infoset from the initial state $w^{0}$, so this backtracking involves at most $|\mathcal{S}_{i}(s_{0})|l^{\text{max}}$ queries to the policy network.
@@
In order to compute the counterfactual values in root infosets for the opponent, let us assume that one of the transformations we used in training is identity transformation, meaning that the critic network output corresponds to the expected value if both players follow the policy from the network, which serves as an approximation of a Nash equilibrium. The opponent’s counterfactual values are the expected values multiplied by the reaches of both chance and resolving players. As discussed, these ranges are always available, so the counterfactual values can always be computed from the definition. Thanks to this, we have both counterfactual values $v^{\pi}_{o}$ and $\beta$ available in each decision point to construct a gadget game and perform a safe search.
We compare previously used continual resolving for constructing public belief state with our approach in Section C.1
4.3 Value Function